# Dual-Spaces Invariance as a Universal Criterion for Identifying Multifractal Critical States

### v5

@@ Abstract
In Anderson localization physics, eigenstates of disordered quantum systems are commonly classified as extended, localized, or critical, distinguished by their distinct spatial structures in real space. While critical states are known to exhibit multifractal characteristics, a precise and operational criterion for characterizing critical states remains an open challenge. In this work, we address this challenge by revisiting criticality from a dual-spaces perspective that treats position and momentum representations on equal footing. Building on the Liu–Xia criterion, which characterizes critical states by the simultaneous vanishing of Lyapunov exponents ($\gamma=\gamma_{m}=0$) in both spaces, we show that this dual-spaces characterization captures an essential feature of critical states that is not limited to Lyapunov exponents. In particular, through numerical simulations, we demonstrate that the inverse participation ratio exhibits closely related scaling behavior in position and momentum space for critical states. This position-momentum correspondence clearly distinguishes critical states from extended or localized ones, which instead display a pronounced asymmetry between the two representations. Our results establish a robust and universal framework for precisely characterizing multifractal critical states in disordered quantum systems, and provide practical guidance for their identification in current quantum simulation platforms.
@@ I Introduction
Anderson localization photonic ; Critical0 is a fundamental manifestation of wave interference in disordered media, leading to the suppression of diffusive transport Liu-self-duality ; Li1 ; Li2 . Within such systems, single-particle eigenstates are commonly classified as extended (metallic), localized (insulating), or critical. Critical states–also referred to as multifractal states–occupy an intermediate regime between extended and localized phases You ; g1 ; zhou . Unlike extended states, which exhibit nearly uniform probability distributions, or localized states, which decay exponentially, critical states typically display power-law spatial decay Biddle ; Zhang ; q2 . Their hallmark features include multifractality, characterized by a continuous spectrum of anomalous scaling exponents q1 ; q4 ; gg10 ; gg11 , as well as scale invariance and statistical self-similarity arising from the competition between localization and delocalization gg12 ; q5 ; q6 . In paradigmatic settings such as the three-dimensional Anderson model, critical states emerge at mobility edges separating extended and localized spectral regions Yao1 ; Yao2 .
@@
Quasiperiodic systems cai ; Slager ; SSHQ ; gg3 ; Ayan1 ; Ayan2 provide an alternative platform for realizing critical states, even in the absence of true randomnes. Recent progress in critical states has uncovered a variety of nontrivial localization phenomena, including anomalous mobility edges Liu-Anomalous , renormalization-group descriptions RG1 ; RG2 , recursion-defined real spectra Liu-Real1 ; Liu-Real2 , and coupled-chain-induced criticality twochain . In previous theoretical descriptions Jiang1 ; Jiang2 ; gg5 ; gg6 , the asymptotic localization properties of single-particle eigenstates were typically characterized by a single real-space Lyapunov exponent $\gamma$. For exponentially localized states, the wave-function amplitudes decay as $|\psi_{n}|\propto e^{-\gamma|n-n_{0}|}$, where $n_{0}$ denotes the localization center and $\gamma>0$ quantifies the inverse localization length gg9 ; Duncan . A finite Lyapunov exponent $\gamma>0$ signals exponential localization, whereas $\gamma=0$ indicates the absence of exponential decay. It is important to emphasize, however, that a vanishing Lyapunov exponent $\gamma=0$ in real space does not uniquely characterize extended states: both truly extended states and critical (multifractal) states fall into the non-exponentially localized class. Consequently, a single-space description based solely on the real-space Lyapunov exponent suffers from intrinsic limitations when applied to critical states Liu-Anomalous .
@@
Figure 1: (Color online) Schematic illustration of typical eigenstates in position and momentum space. Panels (a), (c), and (e) show extended, critical, and localized states in position space, respectively, while panels (b), (d), and (f) show their corresponding momentum-space distributions. In contrast to conventional approaches that emphasize real-space multifractality, a dual-space perspective highlights that critical states are distinguished by the absence of exponential localization in both representations.
@@
A useful conceptual perspective is offered by Bohr’s complementarity principle MQM . Position (real) and momentum spaces constitute a pair of Fourier-conjugate representations: while position space captures the spatial distribution of particles, momentum space encodes their reciprocal structure. Extended and localized states typically display strongly asymmetric localization properties in these two representations. This observation naturally raises the question of whether critical states, which are self-similar by construction, may exhibit a more symmetric behavior under position-momentum transformation.
@@
Motivated by this consideration, Liu and Xia proposed a dual-spaces criterion for the exact identification of critical states Liu-Critical , which overcomes the intrinsic limitations of a single-space description. Within this framework, critical states are uniquely identified by vanishing Lyapunov exponents in both position and momentum space, exhibiting a multifractal yet nonergodic form of delocalization distinct from conventional extended states. By contrast, extended and localized states necessarily exhibit an intrinsic asymmetry between the two conjugate spaces: when the Lyapunov exponent vanishes in one representation, it remains finite in the dual space, as illustrated in Fig. 1(a), (b), (e), and (f). Critical states, on the other hand, are uniquely characterized by the absence of exponential localization in both position and momentum space, as shown in Fig. 1(c) and (d).
@@
The concept of dual-spaces invariance of critical states was first introduced by Liu and Xia in February 2023 on arXiv Liu-Critical , providing a compact and physically transparent framework for classifying eigenstates in Anderson localization. Its applicability has since been supported in a variety of systems, including flat-band lattices quasi1 and Fibonacci topological insulators quasi2 . In March 2025, work quasi3 claimed to characterize critical states through incommensurate zeros of wave functions under dual transformation, following the dual-spaces conceptual framework first introduced by Liu and Xia. These developments underscore the growing interest in dual-spaces approaches to criticality and highlight the foundational role of the dual-spaces invariance framework originally established by Liu et al., which predates and provides the conceptual basis for these subsequent studies.
A key advantage of the Liu–Xia criterion is that it allows for an exact identification of critical states, rather than relying on purely numerical multifractal analysis. Within the multifractal framework, different wave functions are characterized by distinct regimes of the scaling index $\Gamma$ You : in the thermodynamic limit, numerical results for the minimal scaling index quasi4 show that extended states satisfy $\Gamma_{\min}\to 1$, localized states obey $\Gamma_{\min}\to 0$, whereas critical states are characterized by intermediate values $0<\Gamma_{\min}<1$.
By employing the transfer-matrix method (see Appendix A) to compute the Lyapunov exponents $\gamma$ and $\gamma_{m}$ in position and momentum space, respectively, critical states can be unambiguously identified by the simultaneous vanishing of both exponents,
@@
where $E$ denotes the eigenenergy and $V$ the modulation strength. This compact criterion yields an exact characterization of critical states, whose multifractal nature is not readily captured by conventional single-parameter descriptions, as $\Gamma_{\min}$ does not assume a unique value but instead spans the continuous interval $(0,1)$. A unified framework for classifying Anderson localization physics is summarized in Table 1.
Table 1: Classification of typical eigenstates in disordered quantum systems based on the Liu–Xia dual-spaces criterion Liu-Critical .
@@
We stress that the dual-spaces invariance discussed here refers to the invariance of localization characteristics at the level of scaling behavior and statistical properties, rather than an exact equality of critical wave functions. In systems with self-dual symmetry, the critical states at the self-dual point remain invariant under the position-momentum transformation. In contrast, without self-dual symmetry, a critical state in position space is mapped onto a distinct critical state in momentum space, while preserving the same scaling and statistical characteristics. This demonstrates that the dual-spaces invariance implied by the Liu–Xia criterion Liu-Critical is a genuinely nontrivial result. Rather than arising from self-duality that relates identical critical states, it reveals universal features shared by distinct critical states in position and momentum spaces, establishing a fundamentally new understanding of criticality in Anderson localization.
A natural question then arises: beyond Lyapunov exponents, does the dual-space characterization of critical states extend to other physical observables? In this work, we demonstrate that additional experimentally accessible quantities, most notably the inverse participation ratio, exhibit comparable scaling behavior in position and momentum space for critical states. Rather than an exact invariance, this correspondence should be understood at the level of scaling properties and statistical trends. Our results therefore complement the Liu–Xia criterion and provide a practical route for identifying multifractal critical states in both theoretical and experimental settings.
@@ II Phenomenological Analysis
Beyond Lyapunov exponents Longhi1 ; Xia ; Jitomirskaya ; Avila , the inverse participation ratio (IPR) provide complementary information on wave-function localization IPR ; IPR1 ; IPR2 ; IPR3 . The IPR is defined as
@@
for normalized wave functions, and quantifies the degree of spatial confinement of the probability density. Extended states exhibit $\mathrm{IPR}\sim L^{-1}$ in the thermodynamic limit, whereas exponentially localized states are characterized by a finite IPR. Closely related quantities, such as the participation entropy, have also been widely employed as probes of localization and multifractality in both theoretical and experimental studies Fan1 ; Fan2 ; Optical1 ; Optical2 ; gg0 ; gg1 ; gg2 .
For exponentially localized states, the finiteness of the IPR is consistent with the presence of a nonzero Lyapunov exponent. However, for critical states the situation is fundamentally different: their wave functions exhibit power-law decay and strong multifractal fluctuations, which preclude a description in terms of a single characteristic localization length. Consequently, no exact analytical relation between the IPR and the Lyapunov exponent can be generally established for critical states. Nevertheless, from a statistical perspective, the IPR is expected to exhibit behavior analogous to that of the Lyapunov exponent.
Motivated by the Liu–Xia criterion Liu-Critical based on the Lyapunov exponent, we now investigate how the localization properties encoded in the IPR transform between position and momentum spaces. Although the IPR is inherently basis dependent and therefore not invariant under Fourier transformation (see the Appendix B), the dual-spaces invariance underlying the Liu–Xia criterion naturally suggests that critical states should exhibit the same scaling tendency in the two representations. This expectation can be formulated in an operational form as
@@
By contrast, extended and exponentially localized states generically exhibit a pronounced asymmetry between position and momentum spaces, reflected in markedly different scaling behavior of $\mathrm{IPR}$ and $\mathrm{IPR}_{m}$. Critical states, in comparison, are distinguished by a reduced dual-space asymmetry and comparable localization characteristics in both representations, consistent with their multifractal and self-similar nature. On this basis, Table 2 summarizes a phenomenological classification of eigenstates in disordered quantum systems.
@@ III.1 Aubry-André-Harper model
where $\psi_{n}$ is the wave function amplitude of the particle at site $n$, $V_{n}=V\cos(2\pi\alpha n+\theta)$ is the quasiperiodic potential, and $E$ is the energy eigenvalue of the particle. $\alpha$ is an irrational number, usually taken to be the golden ratio $\alpha=\frac{{\sqrt{5}-1}}{2}$, which is a typical choice for quasiperiodicity. $\theta$ is a phase factor that can be adjusted but does not affect the general results (typically set to zero for simplicity).
The AAH model exhibits a remarkable quantum phase transition driven by the quasiperiodic potential strength $V$. In the weak disorder regime ($V<2$), the system maintains extended Bloch-like eigenstates that span the entire lattice, giving rise to metallic transport properties characterized by particle delocalization. Conversely, above the critical threshold ($V>2$), the strong quasiperiodic potential induces Anderson localization, where all eigenstates become exponentially confined in space, completely suppressing particle diffusion. The transition at the critical potential strength $V=2$ represents a unique self-dual point where the system exhibits scale-invariant eigenstates with multifractal characteristics.
@@
Thus, based on Avila’s global theory Avila1 ; Avila2 , the Lyapunov exponent in position space for the AAH model is given by $\gamma=\ln(V/2)$, while in momentum space it is $\gamma_{m}=\ln(2/V)$. Applying the Liu–Xia criterion by setting $\gamma=\gamma_{m}$, we find that the multifractal critical states occur precisely at the potential strength $V=2$. This result is in exact agreement with that obtained from the earlier self-duality relation.
@@ III.2 Quasiperiodic-Nonlinear-Eigenproblem
Non-Hermiticity in the QNE model arises through a mechanism distinct from the three well-established paradigms NH1 ; NH3 ; NH4 : (i) Non-reciprocal hopping (asymmetric tunneling amplitudes), (ii) Complex momentum (gain/loss in reciprocal hopping terms), (iii) Complex on-site potential (gain/loss localized at lattice sites). Instead, the QNE model derives its non-Hermitian character from nonlinear eigenvalue terms, representing an alternative mechanism for generating non-Hermitian spectra, distinct from the conventional paradigms based on asymmetric hopping, gain/loss terms, or complex onsite potentials. This unconventional structure leads to an enlarged critical regime: unlike the AAH model - where critical states appear only at the self-dual point $V=2$ - the QNE model sustains critical states across a broad parameter range $0<V\leq 2$.
After Fourier transformation, see the Appendix C for detailed steps, the real-space QNE Hamiltonian admits exact diagonalization as a linear eigenvalue problem, dramatically reducing computational cost:
@@
Figure 2: (Color online) Panels (a1) and (a2) depict the wave function of the AAH model (with $V=2$, $E=0.0009$) in real and momentum spaces, respectively. Panels (b1) and (b2) display the analogous results for the QNE model at $V=1$, $E=0.4$. The system size is fixed at $L=987$ lattice sites. Notably, the AAH model exhibits self-dual symmetry, under which the critical states remain invariant between position and momentum spaces. By contrast, the QNE model does not possess self-dual symmetry; as a result, a critical state in position space is mapped onto a distinct critical state in momentum space.
We emphasize that the nonlinearity manifests itself in the generalized eigenvalue structure in momentum space, while the real-space representation admits a linear non-Hermitian Hamiltonian. With the explicit forms of the Hamiltonian Eq. (7) and Eq. (8) in both real and momentum space, we can quantitatively study the characteristics of the wave functions in these two representations for the same eigenvalue.

### v6

# Dual-Space Invariance as a Universal Criterion for Multifractal Critical States

@@ Abstract
In Anderson localization, eigenstates of disordered quantum systems are broadly classified as extended, localized, or critical. Although critical states exhibit multifractal character, a precise and operational criterion for their identification remains an open challenge, as Lyapunov exponents in real space cannot uniquely distinguish them from extended states. Here we address this challenge by asserting that critical states are uniquely characterized by an emergent dual-space invariance between position and momentum space. Building on the Liu–Xia criterion of the simultaneous vanishing of Lyapunov exponents ($\gamma=\gamma_{m}=0$), we show that this dual-space invariance extends beyond Lyapunov exponents and governs wavefunction scaling, revealing a fundamental property inaccessible from either space alone. Through numerical simulations, we demonstrate that the inverse participation ratio exhibits matching scaling behavior in position and momentum space for critical states, in sharp contrast to extended and localized states, which display a pronounced asymmetry between the two spaces. This dual-space invariance provides a direct, robust, and universal criterion for identifying multifractal critical states. Our results establish a fundamental principle of Anderson criticality and open new avenues for its detection in modern quantum simulation platforms.
@@ I Introduction
Anderson localization photonic ; Critical0 is a fundamental manifestation of wave interference in disordered media, leading to the suppression of diffusive transport Liu-self-duality ; Li1 ; Li2 . Within such systems, single-particle eigenstates are commonly classified as extended (metallic), localized (insulating), or critical. Critical states–also referred to as multifractal states–occupy an intermediate regime between extended and localized phases You ; g1 ; zhou . Unlike extended states, which exhibit nearly uniform probability distributions, or localized states, which decay exponentially, critical states typically display power-law spatial decay Zhang ; q2 . Their hallmark features include multifractality, characterized by a continuous spectrum of anomalous scaling exponents q1 ; q4 ; gg10 ; gg11 , as well as scale invariance and statistical self-similarity arising from the competition between localization and delocalization gg12 ; q5 ; q6 . In paradigmatic settings such as the three-dimensional Anderson model, critical states occur at mobility edges separating extended and localized regimes NH3 ; Yao1 ; Yao2 .
@@
Quasiperiodic systems cai ; Slager ; SSHQ ; gg3 ; Ayan1 ; Ayan2 provide an alternative platform for realizing critical states, even in the absence of true randomnes. Recent progress in critical states has uncovered a variety of nontrivial localization phenomena, including anomalous mobility edges Liu-Anomalous , renormalization-group descriptions RG1 ; RG2 , recursion-defined real spectra Liu-Real1 ; Liu-Real2 , and coupled-chain-induced criticality twochain . In previous theoretical descriptions Jiang1 ; Jiang2 ; gg5 ; gg6 , the asymptotic localization properties of single-particle eigenstates were typically characterized by a single real-space Lyapunov exponent $\gamma$. The Lyapunov exponent, as a dynamical indicator, signals the localization-delocalization transition. For exponentially localized states, the wave-function amplitudes decay as $|\psi_{n}|\propto e^{-\gamma|n-n_{0}|}$, where $n_{0}$ denotes the localization center and $\gamma>0$ quantifies the inverse localization length gg9 ; Duncan . A finite Lyapunov exponent $\gamma>0$ signals exponential localization, whereas $\gamma=0$ indicates delocalization. It is important to emphasize, however, that a vanishing Lyapunov exponent $\gamma=0$ in real space does not uniquely characterize extended states: both truly extended states and critical (multifractal) states fall into the non-exponentially localized class and exhibit delocalization. Consequently, a single-space description based solely on the real-space Lyapunov exponent suffers from intrinsic limitations when applied to critical states Liu-Anomalous .
@@
Figure 1: (Color online) Schematic illustration of typical eigenstates in position and momentum space. Panels (a), (c), and (e) show extended, critical, and localized states in position space, respectively, while panels (b), (d), and (f) display the corresponding momentum-space distributions. From the dual-space perspective, critical states are uniquely distinguished by the absence of exponential localization in both spaces. This dual absence establishes the Lyapunov exponent as a duality invariant of multifractal critical states.
@@
A useful conceptual perspective is provided by the Fourier uncertainty principle MQM , which dictates that localization in one representation necessarily implies delocalization in its Fourier-dual counterpart. Consequently, localized and extended states exhibit strongly asymmetric localization properties in position (real) and momentum spaces. This naturally raises the question of whether critical states, owing to their self-similar nature, may display a more symmetric behavior under position-momentum transformation.
@@
Motivated by this consideration, Liu and Xia proposed a dual-space criterion for the exact identification of critical states Liu-Critical , overcoming the intrinsic limitations of single-representation descriptions. Specifically, they proposed that critical states lack exponential localization in both position and momentum space, exhibiting vanishing Lyapunov exponents and multifractal, nonergodic delocalization. By contrast, extended and localized states necessarily exhibit an inherent asymmetry between the two conjugate spaces: delocalization in one representation is accompanied by localization in the dual one, as illustrated in Fig. 1(a), (b), (e), and (f). Critical states, however, are uniquely characterized by invariant Lyapunov exponents, indicating delocalization in both position and momentum space, as shown in Fig. 1(c) and (d). This invariance of critical states under the Fourier transform, or dual-space transformation, is termed dual-space invariance.
@@
Dual-space invariance of critical states was first introduced by Liu and Xia in an arXiv preprint in February 2023 Liu-Critical . Its applicability has since been supported in a variety of systems, including flat-band lattices quasi1 and Fibonacci topological insulators quasi2 . In March 2025, Ref. quasi3 claimed to characterize critical states through generalized incommensurate zeros in real and dual spaces, within the dual-space conceptual framework first proposed by Liu and Xia Liu-Critical . These developments underscore the growing interest in dual-space approaches to criticality and further highlight the foundational role of the dual-space invariance framework introduced by Liu et al., which predates and provides the conceptual basis for these subsequent studies.
The notion that critical states are delocalized in both position and momentum space provides an intuitive but oversimplified physical picture. This physical intuition is placed on a rigorous footing by the Liu–Xia criterion, which admits a precise mathematical formulation via the transfer-matrix method (see Appendix A). It identifies critical states through the simultaneous vanishing of the Lyapunov exponents in the two dual spaces,
@@
where $\gamma$ and $\gamma_{m}$ correspond to position and momentum space, respectively, $E$ is the eigenenergy, and $V$ the modulation strength. This implies that the Lyapunov exponent serves as a duality invariant of critical states, and its advantages become evident in comparison with finite-size multifractal analysis.
According to multifractal theory You ; quasi4 , in the thermodynamic limit, the minimal scaling index $\Gamma_{\min}$ approaches $1$ for extended states and 0 for localized states, while critical states satisfy $0<\Gamma_{\min}<1$. However, finite-size multifractal analysis suffers from intrinsic ambiguity, as critical states with small scaling exponents are difficult to distinguish from localized states. The Liu–Xia criterion overcomes this fundamental limitation by providing an exact and representation-independent identification of critical states. This is essential because the multifractal nature of critical states cannot be uniquely characterized by a single scaling exponent, as $\Gamma_{\min}$ spans a continuous interval $(0,1)$ rather than converging to a definite value. Finally, a unified framework for Anderson localization is summarized in Table 1.
Table 1: Classification of typical eigenstates in disordered quantum systems based on the Liu–Xia dual-space criterion Liu-Critical .
@@
A natural question then arises: beyond Lyapunov exponents, does the dual-space characterization of critical states extend to other observables? Here, we show that experimentally accessible quantities, particularly the inverse participation ratio, exhibit analogous scaling in position and momentum space at criticality. This correspondence arises at the level of scaling and statistical properties rather than exact invariance. Our results therefore complement the Liu–Xia criterion and provide a practical route to identifying multifractal critical states in both theoretical and experimental settings.
@@ II Phenomenological Analysis
Compared with the Lyapunov exponent, the inverse participation ratio (IPR) provides complementary information on wave-function localization Longhi1 ; Xia ; Jitomirskaya ; Avila ; IPR ; IPR1 ; IPR2 ; IPR3 . It is defined for normalized wave functions as
@@
and quantifies the spatial confinement of the probability density. Extended states exhibit $\mathrm{IPR}\sim L^{-1}$ in the thermodynamic limit, whereas exponentially localized states are characterized by a finite IPR. Related quantities, such as the participation entropy, have also been widely used to probe localization and multifractality in theoretical and experimental studies Fan1 ; Fan2 ; Optical1 ; Optical2 ; gg0 ; gg1 ; gg2 .
For exponentially localized states, the finite IPR is consistent with a nonzero Lyapunov exponent. In contrast, critical states exhibit power-law decay and strong multifractal fluctuations, which preclude a description in terms of a single localization length. As a result, no exact analytical relation between the IPR and the Lyapunov exponent generally exists for critical states. Nevertheless, from a statistical perspective, the IPR is expected to display behavior analogous to that of the Lyapunov exponent.
Motivated by the Liu–Xia criterion Liu-Critical based on the Lyapunov exponent, we now investigate how the localization properties encoded in the IPR transform between position and momentum spaces. Although the IPR is inherently basis dependent and therefore not invariant under Fourier transformation (see the Appendix B), the dual-space invariance underlying the Liu–Xia criterion naturally suggests that critical states should exhibit the same scaling tendency in the two representations. This expectation can be formulated in an operational form as
@@
By contrast, extended and exponentially localized states typically exhibit a pronounced asymmetry between position and momentum spaces, reflected in the markedly different scaling behavior of $\mathrm{IPR}$ and $\mathrm{IPR}_{m}$. In comparison, critical states remain protected by dual-space symmetry and exhibit comparable localization characteristics in both representations, consistent with their multifractal and self-similar nature. On this basis, Table 2 summarizes a phenomenological classification of eigenstates in disordered quantum systems.
@@
This demonstrates that the Lyapunov exponent itself is an exact duality invariant for multifractal critical states, whereas the IPR, owing to its intrinsic basis dependence, is not strictly duality-invariant. It should be noted that the dual-space invariance discussed here refers to the invariance of localization characteristics at the level of scaling behavior and statistical properties, rather than an exact equality of critical wave functions. In systems with self-dual symmetry, the critical states at the self-dual point remain invariant under the position-momentum transformation. In contrast, without self-dual symmetry, a critical state in position space is mapped onto a distinct critical state in momentum space, while preserving the same scaling and statistical characteristics. This demonstrates that the dual-space invariance implied by the Liu–Xia criterion Liu-Critical is a genuinely nontrivial result. Rather than arising from self-duality that relates identical critical states, it reveals universal features shared by distinct critical states in position and momentum spaces, establishing a fundamentally new understanding of criticality in Anderson localization.
@@ III.1 Aubry-André-Harper model
where $\psi_{n}$ is the wave function amplitude of the particle at site $n$, $V_{n}=V\cos(2\pi\alpha n+\theta)$ is the quasiperiodic potential, and $E$ is the eigenvalue. $\alpha$ is an irrational number, usually taken to be the golden ratio $\alpha=\frac{{\sqrt{5}-1}}{2}$. $\theta$ is a phase factor, typically taken to be zero.
@@
The AAH model exhibits a quantum phase transition driven by the quasiperiodic potential strength $V$. For weak quasiperiodic potential ($V<2$), all eigenstates are extended Bloch-like states, resulting in metallic transport. In contrast, for $V>2$, the system undergoes Anderson localization, where all eigenstates become exponentially localized, suppressing particle diffusion. At the critical point $V=2$, the Hamiltonians in real space [Eq. (4)] and momentum space [Eq. (5)] take identical forms. Consequently, the system reaches the self-dual point and exhibits multifractal eigenstates.
The procedure for identifying multifractal critical states according to the Liu–Xia criterion is outlined as follows. From Avila’s global theory Avila1 ; Avila2 , the Lyapunov exponents in position and momentum spaces are given by $\gamma=\ln(V/2)$ and $\gamma_{m}=\ln(2/V)$, respectively. Imposing the Liu–Xia criterion, $\gamma=\gamma_{m}$, rigorously and uniquely determines the critical point at $V=2$, thereby establishing the multifractal nature of the eigenstates. This result is in exact agreement with that obtained from the self-duality relation, further confirming the validity and consistency of the criterion.
@@ III.2 Quasiperiodic-Nonlinear-Eigenproblem
Non-Hermiticity in the QNE model originates from nonlinear eigenvalue terms, fundamentally distinct from the three established paradigms of asymmetric hopping, gain/loss, and complex potentials NH1 ; NH4 . This unconventional mechanism provides a new route to non-Hermitian spectra. As a direct consequence, the QNE model exhibits an expanded critical regime. Compared with the AAH model, where critical states exist only at the self-dual point $V=2$, the QNE model supports critical states over a broad parameter range $0<V\leq 2$.
After Fourier transformation, the real-space QNE Hamiltonian reduces to a linear eigenvalue problem that can be exactly diagonalized, dramatically reducing the computational cost (see Appendix C for details),
@@
Figure 2: (Color online) Panels (a1) and (a2) show the wave function of the AAH model ($V=2$, $E=0.0009$) in real and momentum space, respectively, while (b1) and (b2) show the corresponding results for the QNE model ($V=1$, $E=0.4$). The system size is $L=987$. Due to self-duality, critical states of the AAH model are invariant between real and momentum space. In contrast, the QNE model lacks self-duality: a critical state transforms into a distinct critical state under Fourier transformation, while their Lyapunov exponents remain identical.
Clearly, unlike the AAH model, the QNE model lacks self-duality symmetry, and its critical states generally exhibit different structures in real and momentum spaces. Consequently, the self-duality relation completely fails to identify the critical states. In sharp contrast, the Liu–Xia criterion remains fully operative. Specifically, the Lyapunov exponents in real and momentum spaces can be rigorously obtained from Eq. (8) and Eq. (7), respectively, and the critical states are uniquely determined by enforcing their equality, as demonstrated in Ref. Liu-Critical . Despite the absence of any self-duality symmetry, the critical states are still dictated by the invariance of the Lyapunov exponents in dual spaces, thereby revealing the fundamental origin, conceptual advance, and universal validity of the Liu–Xia criterion.